IIT JEE Entrance Exam

a stone is dropped from rest passes the top and bottom of a pole at t1 and t2 respectively then time of passing the mid points

zidane dsouza
11 Years agoGrade

1 Answer

Ravi
11 Years ago
Let the top be at height h1 and bottom h2.
So,
under the effect of gravity, the initial velovity being zero,
h1= 1/2(gt12) ...(i)
and h2= 1/2(gt22) ...(ii)
The mid point is at height of h= (h1+h2)/2 which will be equal to 1/2(gt2) ...(iii)

Find the value of t using i, ii and iii